# Dejavu: Towards Experience Feedback Learning for Embodied Intelligence

### v2

@@ B.4.3 R2A-style retrieval-augmented baseline
We also implement a retrieval-augmented RL baseline inspired by R2A [r2a] on top of the same frozen VLA backbone as EFN. Because the original R2A is designed for value-based, non-residual control with a specialized dual-process architecture and multiple auxiliary objectives, we do not attempt a bit-level reproduction. Instead, we construct an R2A-style retrieval-augmented SAC that preserves the core idea of conditioning the policy and critic on a retrieval-derived context vector, while adapting the design to our continuous-control, residual-on-VLA setting. Throughout this section we explicitly refer to this baseline as an adapted implementation and keep all SAC hyper-parameters and backbone interfaces identical to those of EFN for a fair comparison.
@@
$s_{t}=\big[f_{\text{vis}}(o_{t}),\,f_{\text{prop}}(x_{t}),\,e(\ell)\big],$ (38)
@@
$\alpha_{j}=\operatorname{softmax}_{j}\!\left(\frac{q_{t}^{\top}W_{a}\tilde{s}_{i_{j}}}{\sqrt{d}}\right),\qquad u_{t}=\sum_{j=1}^{K}\alpha_{j}\,W_{v}\tilde{s}_{i_{j}},$ (39)
@@ B.4.4 GC-TTT-style test-time training baseline
$\mathcal{L}_{\text{GC-TTT}}(\phi)=-\mathbb{E}_{(s_{i},a_{i})\sim\mathcal{D}_{\text{good}}}\big[\log\pi_{\phi}(a_{i}\mid s_{i},g^{\ast})\big],$ (41)
@@ C.2 Why a Frozen VLA with a Residual Head?
$a_{t}=a_{t}^{\text{base}}+\Delta a_{t}.$ (42)
This preserves the backbone’s strong generalization and task coverage, while allowing EFN to refine local decisions where the base policy systematically underperforms. Because the residual only needs to model deviations from a competent policy, it is more sample-efficient and typically more stable to train, consistent with the intuition behind residual policy learning and ResAct-style methods [resact].
Our ablations further support this design. When we replace the residual head with a standalone action predictor that ignores the base action (“no residual / direct action prediction”), performance drops in both success rate and step efficiency (see Appendix E). In other words, throwing away the base action and asking the small head to solve the entire control problem again is strictly worse than learning corrections on top of it.
@@ C.3 Why an Episodic Experience Bank?
$\mu$ $=\frac{1}{N}\sum_{i}v_{i},$ (43)
$m$ $=\max_{i}v_{i}\quad(\text{element-wise}),$ (44)
$k$ $=\operatorname{LN}([\mu;m]),$ (45)
@@
We further impose two simple but important priors in retrieval: instruction filtering and an efficiency prior. Instruction filtering uses $\ell_{\tau}$ to restrict retrieval to episodes whose language embedding is close to the current instruction, which reduces harmful cross-task transfer on multi-task benchmarks such as LIBERO and prevents the agent from imitating “good actions for the wrong task.” The efficiency prior biases the bank towards shorter successful trajectories: during offline construction and online retrieval, we prefer episodes that reach success in fewer steps. This aligns with deployment-time goals (reusing concise, high-quality behaviors) and empirically improves both success and step efficiency; removing instruction filtering or the length prior (“w/o instruction embed”) degrades performance in our ablations (see Appendix Table 6).
@@
(b) Calvin tasks
(c) Real-world tasks
Figure 9: Performance consistently improves as the experience bank grows: (a) simulated LIBERO tasks with OpenVLA, (b) simulated CALVIN tasks with OpenVLA, and (c) real-world manipulation tasks with GO-1. Here, the bank volume refers to the number of stored trajectories (episodes) in the experience bank.
@@ C.4 Why Residual Policy Optimization Instead of Direct Action Prediction?
$a_{t}\;=\;a_{t}^{\text{base}}+\Delta a_{t}.$ (46)
@@ C.5 Why Similarity-shaped Dense Rewards and the Anti-idling Design?
The reward decomposition in Figure 8 is designed to turn retrieval into a dense and behaviorally aligned learning signal, rather than relying solely on sparse task success. A naive starting point is to reward the agent by how well its next observation matches the retrieved next frame, e.g., a cosine similarity term $a_{t}\in[0,1]$ between the actual next visual embedding and the retrieved target. This already provides a dense signal at every step, but it also introduces a failure mode: the agent can exploit the reward by finding a visually similar pose and then “vibrating in place,” maintaining high similarity without making real progress towards task completion.
@@
From this perspective, the reward is not a black-box signal delegated entirely to SAC to “figure out” credit assignment; instead, it directly encodes our behavioral preferences: (i) match good retrieved futures, (ii) move towards them rather than hovering idly nearby, and (iii) reach them quickly. The ablations in Table 6 (also summarized in Figure 8) support this design. Removing the similarity-shaped dense terms (“w/o dense rewards”) and relying only on sparse task success leads to lower average success and less consistent improvements over OpenVLA, indicating that the dense retrieval-based components are crucial for effective post-deployment learning. Dropping the anti-idle terms (“w/o anti-idle”) yields a policy that attains similar success but with noticeably longer episodes, confirming that, without explicit penalties on idleness, the agent tends to favor visually safe but inefficient behaviors. In contrast, the full EFN reward decomposes the similarity signal into progress, motion, anti-idle, and time components, producing the best overall trade-off between success and step efficiency on LIBERO.
@@ C.6 Why Reinforcement Learning for EFN?
$r^{\text{sem}}_{t}\;=\;\cos\!\big(\mathbf{u}(\mathbf{F}_{t+1}),\,\mathbf{u}(\hat{\mathbf{F}}^{+})\big),$ (47)
@@
Our ablations support these design decisions. When we remove SAC and rely only on a value-style critic without the entropy-regularized actor update (“w/o SAC” in Table 6), EFN’s improvements over the backbone shrink in both success and step metrics. In contrast, the full SAC-based optimization leverages the dense similarity rewards and the retrieval structure to deliver consistent gains on top of a frozen VLA policy. Overall, reinforcement learning—and SAC in particular—fits our design philosophy: a lightweight optimization layer that operates on the VLA’s latent interface, reuses interaction experience, and learns directly from deployment-time feedback without modifying any backbone weights.
Figure 10: Composition of successful, near-successful, and failed trajectories on Libero.
Figure 11: Per-step reward distribution for successful, near-successful, and failed episodes.
Figure 12: Training performance with different experience bank compositions on Libero.
Figure 13: Comparison of rollout update strategies during deployment: Success-only vs All.
@@ C.7 Why Live Experience Growth at Deployment?
Freezing weights does not preclude “learning from experience”: EFN adapts by changing which trajectories it can retrieve, not how it processes them. As shown in Figure 9, enlarging the experience bank consistently improves success rates on LIBERO, CALVIN, and real-world GO-1 tasks, because a larger bank provides more relevant and diverse execution snippets for residual guidance on top of the frozen VLA. By appending only validated (successful) rollouts, this live growth remains a safe form of adaptation: the agent gains from accumulating high-quality experiences while preserving the stability, compute efficiency, and auditability of a frozen model.
@@ D.1 Success and Failure Experience
A key design choice in EFN is to treat success, near-success, and failure differently at training and deployment time. During training, the offline bank intentionally contains a mixture of successful rollouts, near-successful rollouts that complete only a subset of sub-goals, and even clearly failed trajectories (Figure  13). This is feasible because our reward $r_{t}$ is defined purely from semantic feature similarity rather than a binary success flag: any transition that moves the state closer to a goal-consistent embedding can provide useful signal, regardless of whether the episode ultimately succeeds.
Figures 13 and 13 support this design. The per-step reward distributions show that failed and near-successful episodes still contain many high-similarity steps, i.e., meaningful partial progress. In the ablation on bank composition, moving from S-only to S+NS consistently improves performance, and even S+NS+F does not significantly degrade results, indicating that EFN is robust to a moderate amount of failure noise when the reward is similarity-based.
In contrast, during deployment we only append rollouts classified as successful to the live bank. Success detection is fully automatic: instead of human labels, we reuse the same feature-matching criterion as Eq. (11) in the main paper and mark an episode as successful if its final state embedding is sufficiently similar to the goal embedding. Figure 13 shows that appending all rollouts (All-rollouts) gradually harms performance, whereas a success-only update strategy maintains a cleaner bank and yields higher, more stable success rates. A remaining limitation is that our current detector is a simple similarity threshold; future work could explore learned, task-aware success estimators or soft weighting of borderline near-successes to further refine which experiences are admitted into the live bank.
[…]
Method Success@1 Success@2 Success@3 Success@4 Success@5 Avg. Len $\downarrow$ OpenVLA 0.908 0.765 0.612 0.515 0.430 3.230 +kNN-RAG 0.902 0.755 0.601 0.498 0.412 3.168 +ResAct [resact] 0.928 0.793 0.645 0.542 0.458 3.366 +R2A [r2a] 0.936 0.810 0.662 0.557 0.472 3.437 +GC-TTT 0.921 0.785 0.638 0.535 0.448 3.327 +EFN (Vol=2000) 0.945 0.828 0.685 0.578 0.495 3.531 +EFN (Vol=5000) 0.953 0.842 0.701 0.592 0.508 3.596 UniVLA 0.948 0.845 0.741 0.656 0.552 3.742 +kNN-RAG 0.942 0.831 0.726 0.642 0.548 3.689 +ResAct [resact] 0.960 0.865 0.768 0.682 0.579 3.854 +R2A [r2a] 0.968 0.877 0.783 0.696 0.591 3.915 +GC-TTT 0.958 0.862 0.764 0.678 0.575 3.837 +EFN (Vol=2000) 0.973 0.885 0.796 0.712 0.608 3.974 +EFN (Vol=5000) 0.978 0.894 0.808 0.725 0.621 4.026
Table 6: Ablation Studies on Libero Dataset
@@ D.2 Experience-centric post-deployment learning
Viewing post-deployment adaptation as experience growth rather than perpetual finetuning has broader implications. It makes behavior easier to reproduce and audit, since changes over time are driven by the content of an explicit memory store rather than opaque weight drift, and it aligns better with realistic deployment constraints where on-board compute and engineering budgets limit extensive online training. At the same time, it shifts the core challenges toward curating, scaling, and querying experience banks. For example, dealing with long-term storage, retrieval efficiency, and coverage of rare but safety-critical events. A more detailed rationale for each architectural choice underlying this experience-centric design is given in Appendix C; here we emphasize the conceptual shift from parameter-centric to experience-centric post-deployment learning.
@@ D.3 When does EFN help, and why?
Empirically, EFN helps most once the experience bank covers a modest set of typical situations rather than an extremely large corpus. Increasing the bank volume from 0 to $300$ episodes already yields a major jump in performance, while further expanding to $1000$ provides only diminishing but consistent gains (Figure  9). This suggests that EFN is mainly sensitive to whether common layouts, viewpoints, and language instructions are represented, not to exhaustively storing every trajectory. In benchmarks like LIBERO and CALVIN, where tasks share similar manipulation structure, even a medium-sized bank is enough for retrieval to regularly surface useful precedents.
@@ D.4 Limitations
The first limitation of EFN is its reliance on the backbone representation quality. EFN operates entirely in the latent space exposed by the frozen VLA; if the backbone fails catastrophically for certain objects, viewpoints, or language instructions, both retrieval and residual correction will fail jointly. Moreover, the current design needs a non-trivial number of successful trajectories before it can boost the performance: in extremely cold-start regimes with virtually no successes, EFN has little useful signal to retrieve from. In practice, we observe that once there are even a handful of successful rollouts, EFN can already yield substantial gains. For example, on the challenging DrawerStore task in the real-world AgiBot-G1 experiments, whose base success rate is only 5.3%, EFN increases performance to 37.3% with a bank volume of 300 and further to 58.7% with volume 1000 (Table 2; visualizations in Section F). Nevertheless, if the backbone truly never succeeds (e.g., close to 0% success), our method provides limited leverage and the more appropriate remedy is to retrain or replace the underlying VLA rather than rely on post-deployment experience alone.
The second limitation concerns the size and management of the experience bank. Under the scales considered in this work (LIBERO simulation and our AgiBot-G1 experiments), the memory footprint and retrieval latency remain modest, and EFN adds only a small overhead on top of the base VLA policy as summarized in Table 3. However, in long-running deployments with many robots or continuously operating fleets, the bank will grow without bound unless additional mechanisms are introduced. In such settings, it will be necessary to design more sophisticated caching, compression, and replacement strategies (e.g., task-aware coreset selection, time- or novelty-based forgetting, or hierarchical multi-bank structures) to keep both storage and retrieval cost under control while preserving the benefits of rich episodic experience.
@@ Appendix E Supplementary Experimental Results
E.1 Analysis on Calvin Dataset
[…]
Does EFN require nearly identical scenes?
[…]
E.2 Runtime and Memory Footprint of EFN
E.2.1 Standalone Cost
@@
Table 7: Standalone runtime and memory footprint.
@@
We benchmark $200$ forward passes after $50$ warm-up iterations using CUDA event timing. As shown in Table 7, EFN adds only $\sim\!1.94$ ms per control step and incurs minimal activation memory ($\sim\!2.6$ MB). The parameter footprint is modest ($208$ MB), especially compared to the frozen VLA backbone. These results indicate that EFN is lightweight enough for real-time control and can be cleanly integrated into existing policies.
E.2.2 Integration Cost
We next examine the end-to-end overhead when EFN is integrated into the full GO-1 policy, including vision encoding, LLM inference, and action decoding. This measurement in Table 8 reflects the actual burden added during real-world deployment.
Table 8: Inference-time efficiency comparison on real-world deployment. Change is computed relative to the Base GO-1.

### v3

@@ B.4.3 R2A-style retrieval-augmented baseline
We also implement a retrieval-augmented RL baseline inspired by R2A [13] on top of the same frozen VLA backbone as EFN. Because the original R2A is designed for value-based, non-residual control with a specialized dual-process architecture and multiple auxiliary objectives, we do not attempt a bit-level reproduction. Instead, we construct an R2A-style retrieval-augmented SAC that preserves the core idea of conditioning the policy and critic on a retrieval-derived context vector, while adapting the design to our continuous-control, residual-on-VLA setting. Throughout this section we explicitly refer to this baseline as an adapted implementation and keep all SAC hyper-parameters and backbone interfaces identical to those of EFN for a fair comparison.
@@
$s_{t}=\big[f_{\text{vis}}(o_{t}),\,f_{\text{prop}}(x_{t}),\,e(\ell)\big],$ (B.22)
@@
$\alpha_{j}=\operatorname{softmax}_{j}\!\left(\frac{q_{t}^{\top}W_{a}\tilde{s}_{i_{j}}}{\sqrt{d}}\right),\qquad u_{t}=\sum_{j=1}^{K}\alpha_{j}\,W_{v}\tilde{s}_{i_{j}},$ (B.23)
@@ B.4.4 GC-TTT-style test-time training baseline
$\mathcal{L}_{\text{GC-TTT}}(\phi)=-\mathbb{E}_{(s_{i},a_{i})\sim\mathcal{D}_{\text{good}}}\big[\log\pi_{\phi}(a_{i}\mid s_{i},g^{\ast})\big],$ (B.25)
@@ C.2 Why a Frozen VLA with a Residual Head?
$a_{t}=a_{t}^{\text{base}}+\Delta a_{t}.$ (C.1)
This preserves the backbone’s strong generalization and task coverage, while allowing EFN to refine local decisions where the base policy systematically underperforms. Because the residual only needs to model deviations from a competent policy, it is more sample-efficient and typically more stable to train, consistent with the intuition behind residual policy learning and ResAct-style methods [29].
Our ablations further support this design. When we replace the residual head with a standalone action predictor that ignores the base action (“no residual / direct action prediction”), performance drops in both success rate and step efficiency (see Appendix Section E.2). In other words, throwing away the base action and asking the small head to solve the entire control problem again is strictly worse than learning corrections on top of it.
@@ C.3 Why an Episodic Experience Bank?
$\mu$ $=\frac{1}{N}\sum_{i}v_{i},$ (C.2)
$m$ $=\max_{i}v_{i}\quad(\text{element-wise}),$ (C.3)
$k$ $=\operatorname{LN}([\mu;m]),$ (C.4)
@@
We further impose two simple but important priors in retrieval: instruction filtering and an efficiency prior. Instruction filtering uses $\ell_{\tau}$ to restrict retrieval to episodes whose language embedding is close to the current instruction, which reduces harmful cross-task transfer on multi-task benchmarks such as LIBERO and prevents the agent from imitating “good actions for the wrong task.” The efficiency prior biases the bank towards shorter successful trajectories: during offline construction and online retrieval, we prefer episodes that reach success in fewer steps. This aligns with deployment-time goals (reusing concise, high-quality behaviors) and empirically improves both success and step efficiency; removing instruction filtering or the length prior (“w/o instruction embed”) degrades performance in our ablations (see Appendix Table D.1).
@@
(b) Real-world tasks
Figure C.1: Performance consistently improves as the experience bank grows in both simulated LIBERO tasks with OpenVLA and real-world manipulation tasks with GO-1. Here, the bank volume refers to the number of stored trajectories (episodes) in the experience bank.
@@ C.4 Why Residual Policy Optimization Instead of Direct Action Prediction?
$a_{t}\;=\;a_{t}^{\text{base}}+\Delta a_{t}.$ (C.5)
@@ C.5 Why Similarity-shaped Dense Rewards and the Anti-idling Design?
The reward decomposition in main-paper Fig. 4 is designed to turn retrieval into a dense and behaviorally aligned learning signal, rather than relying solely on sparse task success. A naive starting point is to reward the agent by how well its next observation matches the retrieved next frame, e.g., a cosine similarity term $a_{t}\in[0,1]$ between the actual next visual embedding and the retrieved target. This already provides a dense signal at every step, but it also introduces a failure mode: the agent can exploit the reward by finding a visually similar pose and then “vibrating in place,” maintaining high similarity without making real progress toward task completion.
@@
From this perspective, the reward is not a black-box signal delegated entirely to SAC to “figure out” credit assignment; instead, it directly encodes our behavioral preferences: (i) match good retrieved futures, (ii) move toward them rather than hovering idly nearby, and (iii) reach them quickly. The ablations in Table D.1 (also summarized in main-paper Fig. 4) support this design. Removing the similarity-shaped dense terms (“w/o dense rewards”) and relying only on sparse task success leads to lower average success and less consistent improvements over OpenVLA, indicating that the dense retrieval-based components are crucial for effective post-deployment learning. Dropping the anti-idle terms (“w/o anti-idle”) yields a policy that attains similar success but with noticeably longer episodes, confirming that, without explicit penalties on idleness, the agent tends to favor visually safe but inefficient behaviors. In contrast, the full EFN reward decomposes the similarity signal into progress, motion, anti-idle, and time components, producing the best overall trade-off between success and step efficiency on LIBERO.
@@ C.6 Why Reinforcement Learning for EFN?
$r^{\text{sem}}_{t}\;=\;\cos\!\big(\mathbf{u}(\mathbf{F}_{t+1}),\,\mathbf{u}(\hat{\mathbf{F}}^{+})\big),$ (C.6)
@@
Our ablations support these design decisions. When we remove SAC and rely only on a value-style critic without the entropy-regularized actor update (“w/o SAC” in Table D.1), EFN’s improvements over the backbone shrink in both success and step metrics. In contrast, the full SAC-based optimization leverages the dense similarity rewards and the retrieval structure to deliver consistent gains on top of a frozen VLA policy. Overall, reinforcement learning—and SAC in particular—fits our design philosophy: a lightweight optimization layer that operates on the VLA’s latent interface, reuses interaction experience, and learns directly from deployment-time feedback without modifying any backbone weights.
Figure C.2: Composition of successful, near-successful, and failed trajectories on LIBERO.
Figure C.3: Per-step reward distribution for successful, near-successful, and failed episodes.
Figure C.4: Training performance with different experience bank compositions on LIBERO.
Figure C.5: Comparison of rollout update strategies during deployment: Success-only vs All.
@@ C.7 Why Live Experience Growth at Deployment?
Freezing weights does not preclude “learning from experience”: EFN adapts by changing which trajectories it can retrieve, not how it processes them. As shown in Figure C.1, enlarging the experience bank consistently improves success rates on LIBERO and real-world GO-1 tasks, because a larger bank provides more relevant and diverse execution snippets for residual guidance on top of the frozen VLA. By appending only validated (successful) rollouts, this live growth remains a safe form of adaptation: the agent gains from accumulating high-quality experiences while preserving the stability, compute efficiency, and auditability of a frozen model.
@@ D.1 Success and Failure Experience
A key design choice in EFN is to treat success, near-success, and failure differently at training and deployment time. During training, the offline bank intentionally contains a mixture of successful rollouts, near-successful rollouts that complete only a subset of sub-goals, and even clearly failed trajectories (Figure C.5). This is feasible because our reward $r_{t}$ is defined purely from semantic feature similarity rather than a binary success flag: any transition that moves the state closer to a goal-consistent embedding can provide useful signal, regardless of whether the episode ultimately succeeds.
Figures C.5 and C.5 support this design. The per-step reward distributions show that failed and near-successful episodes still contain many high-similarity steps, i.e., meaningful partial progress. In the ablation on bank composition, moving from S-only to S+NS consistently improves performance, and even S+NS+F does not significantly degrade results, indicating that EFN is robust to a moderate amount of failure noise when the reward is similarity-based.
In contrast, during deployment we only append rollouts classified as successful to the live bank. Success detection is fully automatic: instead of human labels, we reuse the same feature-matching criterion as Eq. (11) in the main paper and mark an episode as successful if its final state embedding is sufficiently similar to the goal embedding. Figure C.5 shows that appending all rollouts (All-rollouts) gradually harms performance, whereas a success-only update strategy maintains a cleaner bank and yields higher, more stable success rates. A remaining limitation is that our current detector is a simple similarity threshold; future work could explore learned, task-aware success estimators or soft weighting of borderline near-successes to further refine which experiences are admitted into the live bank.
Table D.1: Ablation studies on the LIBERO benchmark
@@ D.2 Experience-centric post-deployment learning
Viewing post-deployment adaptation as experience growth rather than perpetual finetuning has broader implications. It makes behavior easier to reproduce and audit, since changes over time are driven by the content of an explicit memory store rather than opaque weight drift, and it aligns better with realistic deployment constraints where on-board compute and engineering budgets limit extensive online training. At the same time, it shifts the core challenges toward curating, scaling, and querying experience banks, for example by addressing long-term storage, retrieval efficiency, and coverage of rare but safety-critical events. A more detailed rationale for each architectural choice underlying this experience-centric design is given in Appendix Section C; here we emphasize the conceptual shift from parameter-centric to experience-centric post-deployment learning.
@@ D.3 When does EFN help, and why?
Empirically, EFN helps most once the experience bank covers a modest set of typical situations rather than an extremely large corpus. Increasing the bank volume from 0 to $300$ episodes already yields a major jump in performance, while further expanding to $1000$ provides only diminishing but consistent gains (Figure C.1). This suggests that EFN is mainly sensitive to whether common layouts, viewpoints, and language instructions are represented, rather than to exhaustively storing every trajectory. In benchmarks like LIBERO, where tasks share similar manipulation structure, even a medium-sized bank is enough for retrieval to regularly surface useful precedents.
@@ D.4 Limitations
The first limitation of EFN is its reliance on the backbone representation quality. EFN operates entirely in the latent space exposed by the frozen VLA; if the backbone fails catastrophically for certain objects, viewpoints, or language instructions, both retrieval and residual correction will fail jointly. Moreover, the current design needs a non-trivial number of successful trajectories before it can boost performance: in extremely cold-start regimes with virtually no successes, EFN has little useful signal to retrieve. In practice, we observe that once there are even a handful of successful rollouts, EFN can already yield substantial gains. For example, on the challenging DrawerStore task in the real-world AgiBot-G1 experiments, whose base success rate is only 5.3%, EFN increases performance to 37.3% with a bank volume of 300 and further to 58.7% with volume 1000 (Table 2 in the main paper; visualizations in Section F). Nevertheless, if the backbone truly never succeeds (e.g., close to $0\%$ success), our method provides limited leverage, and the more appropriate remedy is to retrain or replace the underlying VLA rather than rely on post-deployment experience alone.
The second limitation concerns the size and management of the experience bank. Under the scales considered in this work (LIBERO simulation and our AgiBot-G1 experiments), the memory footprint and retrieval latency remain modest, and EFN adds only a small overhead on top of the base VLA policy, as summarized in Table 3 in the main paper. However, in long-running deployments with many robots or continuously operating fleets, the bank will grow without bound unless additional mechanisms are introduced. In such settings, it will be necessary to design more sophisticated caching, compression, and replacement strategies (e.g., task-aware coreset selection, time- or novelty-based forgetting, or hierarchical multi-bank structures) to keep both storage and retrieval cost under control while preserving the benefits of rich episodic experience.
@@ Appendix E Supplementary Experimental Results
E.1 Runtime and Memory Footprint of EFN
E.1.1 Standalone Cost
@@
Table E.1: Standalone runtime and memory footprint.
@@
We benchmark $200$ forward passes after $50$ warm-up iterations using CUDA event timing. As shown in Table E.1, EFN adds only $\sim\!1.94$ ms per control step and incurs minimal activation memory ($\sim\!2.6$ MB). The parameter footprint is modest ($208$ MB), especially compared to the frozen VLA backbone. These results indicate that EFN is lightweight enough for real-time control and can be cleanly integrated into existing policies.
E.1.2 Integration Cost
We next examine the end-to-end overhead when EFN is integrated into the full GO-1 policy, including vision encoding, LLM inference, and action decoding. This measurement in Table E.2 reflects the actual burden added during real-world deployment.
Table E.2: Inference-time efficiency comparison for real-world deployment. Change is computed relative to the base GO-1.